# New mixed inhomogeneous phase in vortical gluon plasma: first-principle results from rotating SU(3) lattice gauge theory

### v2

@@ Introduction.
We claim that the violation of the TE law is a consequence of the conformal anomaly in analogy with a similar effect in condensed matter systems Bermond et al. (2022). We suggest that the negative moment of inertia observed in our previous studies Braguta et al. (2023b, c, d, e) and the violation of the Tolman-Ehrenfest law found in this paper are linked to each other through a yet-to-be-discovered relationship. This link remains to be explored and is left for future work.
Acknowledgements.
The work of VVB and AAR has been carried out using computing resources of the Federal collective usage center Complex for Simulation and Data Processing for Mega-science Facilities at NRC “Kurchatov Institute”, http://ckp.nrcki.ru/ and the Supercomputer “Govorun” of Joint Institute for Nuclear Research. The work of VVB and AAR, which consisted of the lattice calculation of the observables used in the paper and interpretation of the data, was supported by the Russian Science Foundation (project no. 23-12-00072). The work of MNC has been supported by the French National Agency for Research (ANR) within the project PROCURPHY. MNC is thankful to the members of Nordita (Stockholm) for their kind hospitality. Nordita is supported in part by Nordforsk. The authors are grateful to Dmitrii Sychev and Andrey Kotov for useful discussions. MNC thanks Frank Wilczek for valuable comments.

### v3

@@ 11 Summary and Conclusions
We claim that the violation of the TE law is a consequence of the conformal anomaly in analogy with a similar effect in condensed matter systems [46]. We suggest that the negative moment of inertia observed in our previous studies [26, 29, 30, 31] and the violation of the Tolman-Ehrenfest law found in this paper are linked to each other through a yet-to-be-discovered relationship. This link remains to be explored and is left for future work.
Acknowledgements
The work of VVB and AAR has been carried out using computing resources of the Federal collective usage center Complex for Simulation and Data Processing for Mega-science Facilities at NRC “Kurchatov Institute”, http://ckp.nrcki.ru/ and the Supercomputer “Govorun” of Joint Institute for Nuclear Research. The work of VVB and AAR, which consisted of the lattice calculation of the observables used in the paper and interpretation of the data, was supported by the Russian Science Foundation (project no. 23-12-00072). The work of MNC has been supported by the French National Agency for Research (ANR) within the project PROCURPHY ANR-23-CE30-0051-02. MNC is thankful to the members of Nordita (Stockholm) for their kind hospitality. Nordita is supported in part by Nordforsk. The authors are grateful to Dmitrii Sychev and Andrey Kotov for useful discussions. MNC thanks Frank Wilczek for valuable comments.